# A Culturally-diverse Multilingual Multimodal Video Benchmark & Model

### v2

@@ Abstract
Large multimodal models (LMMs) have recently gained attention due to their effectiveness to understand and generate descriptions of visual content. Most existing LMMs are in English language. While few recent works explore multilingual image LMMs, to the best of our knowledge, moving beyond the English language for cultural and linguistic inclusivity is yet to be investigated in the context of video LMMs. In pursuit of more inclusive video LMMs, we introduce a multilingual Video LMM benchmark, named ViMUL-Bench, to evaluate Video LMMs across 14 languages, including both low- and high-resource languages: English, Chinese, Spanish, French, German, Hindi, Arabic, Russian, Bengali, Urdu, Sinhala, Tamil, Swedish, and Japanese. Our ViMUL-Bench is designed to rigorously test video LMMs across 15 categories including eight culturally diverse categories, ranging from lifestyles and festivals to foods and rituals and from local landmarks to prominent cultural personalities. ViMUL-Bench comprises both open-ended (short and long-form) and multiple-choice questions spanning various video durations (short, medium, and long) with 8k samples that are manually verified by native language speakers. In addition, we also introduce a machine translated multilingual video training set comprising 1.2 million samples and develop a simple multilingual video LMM, named ViMUL, that is shown to provide a better tradeoff between high-and low-resource languages for video understanding. We hope our ViMUL-Bench and multilingual video LMM along with a large-scale multilingual video training set will help ease future research in developing cultural and linguistic inclusive multilingual video LMMs. Our proposed benchmark, video LMM codebase and training data are available at https://mbzuai-oryx.github.io/ViMUL/.
@@ A Culturally-diverse Multilingual Multimodal Video Benchmark & Model
Bhuiyan Sanjid Shafique1thanks: Equal contribution, Ashmal Vayani211footnotemark: 1, Muhammad Maaz1, Hanoona Abdul Rasheed1, Dinura Dissanayake1, Mohammed Irfan Kurpath1, Yahya Hmaiti2 Go Inoue1, Jean Lahoud1, Md. Safirur Rashid3, Shadid Intisar Quasem3, Maheen Fatima4 Franco Vidal2, Mykola Maslych2, Ketan Pravin More1, Sanoojan Baliah1 Hasindri Watawana1, Yuhao Li1, Fabian Farestam5, Leon Schaller6, Roman Tymtsiv7 Simon Weber6, Hisham Cholakkal1, Ivan Laptev1, Shin’ichi Satoh8, Michael Felsberg10 Mubarak Shah2, Salman Khan1,9, Fahad Shahbaz Khan1,10 1Mohamed bin Zayed University of AI, 2University of Central Florida, 3Islamic University of Technology, 4Air University 5ETH Zurich, 6 Technische Universität München, 7Independent Researcher, 8National Institute of Informatics 9Australian National University, 10Linköping University {bhuiyan.shafique, muhammad.maaz, hanoona.bangalath, dinura.dissanayake, mohammedirfan.k, go.inoue, jean.lahoud ketan.more, sanoojan.baliah, hasindri.watawana, hisham.cholakkal, salman.khan, fahad.khan}@mbzuai.ac.ae {ashmal.vayani, yohan.hmaiti, fr543419, mykola.maslych}@ucf.edu, {safirurrashid, shadidintisar}@iut-dhaka.edu {leon.schaller, simon1.weber}@tum.de, {tymtsiv.roman, lyh88524}@gmail.com, maheen.fatima@students.au.edu.pk ffarestam@student.ethz.ch, satoh@nii.ac.jp, michael.felsberg@liu.se, ivan.laptev@inria.fr, mubarak.shah@crcv.ucf.edu
@@ 1 Introduction
Existing efforts to explore linguistically and culturally diverse LMM benchmarks are limited to images Romero et al. (2024); Vayani et al. (2025). To the best of our knowledge, linguistic and cultural diversity are yet to be investigated for video LMMs. Video domain poses different challenges as it often depicts complex, culturally rich scenarios-local festivals, foods, rituals, or landmarks, that require understanding both the visual context and the language-specific narration or questions. Fig. 1 illustrates an example where LMMs are asked to describe the taste of the Bengali dish Beef Tehari. The models fail to interpret the question in the local language, responding incorrectly and missing linguistic nuances. While short and long video understanding LMM benchmarks exist in literature, they are typically restricted to only English language. For instance, Video-MME Fu et al. (2024) focuses on diverse video analysis but in a single language, and MVBench Li et al. (2024b) emphasizes temporal reasoning (action sequences, motion) without multilingual considerations. Other recent efforts like ViLMA Kesen et al. (2024) and SEED-Bench Li et al. (2023b, a) probe video-language model’s abilities in zero-shot temporal grounding and procedural understanding, among other skills, yet none assess cross-lingual or cross-cultural comprehension. In short, there is no comprehensive benchmark to evaluate how well video LMMs perform across different languages and cultural contexts (see Tab. 1).
@@
  • We introduce ViMUL-Bench, a comprehensive benchmark for video LMMs covering 14 languages (including several under-represented ones) and 15 domains including real-world cultural aspects. To our knowledge, this is the first effort to enable rigorous testing of video LMMs across a wide linguistic and cultural spectrum, emphasizing both cross-lingual and cultural comprehension (see Tab. 1).
@@ 2 Related Works
Table 1: Comparison of video LMM benchmarks emphasizing multilingual and cultural understanding. Domains represent the aspects covered by each dataset for different languages. Annotation Type is categorized as follows: Human - Questions were created in the local language. Human+Auto - Questions were generated or translated using GPT-4/Google API and later validated by human experts. Auto: Questions were generated or translated automatically without human validation. ’-’ indicates that information is not available.
@@
Video LMM Benchmarks: Extending multimodal evaluation to video introduces additional challenges due to temporal dynamics. A number of video-language benchmarks have emerged to assess LMMs on video understanding, though they focus on general monolingual capabilities. Video-MME Fu et al. (2024) is introduced as the first comprehensive evaluation of LMMs on video analysis, covering a broad spectrum of video domains. It offers a full-spectrum evaluation, covering questions from perception to reasoning, and variable videos lengths ($\sim$10 secs to $\sim$1 hour). Another effort, MVBench Li et al. (2024b) concentrates on temporal reasoning skills and proposes an automatic pipeline to generate a large multiple-choice QA benchmark by leveraging existing video datasets and GPT-based annotators. Beyond general benchmarks, ViLMA Kesen et al. (2024) takes a fine-grained approach as it uses carefully designed counterfactual video pairs to probe a model’s temporal grounding and linguistic understanding in a zero-shot setting. SEED-Bench is another comprehensive multimodal benchmark that includes some video-based questions; however, those are largely confined to temporal or procedural understanding tasks Li et al. (2023b, a). These benchmarks have a significantly advanced evaluation for video-based multimodal reasoning and generation, however they lack multilingual and culturally aware components. Current video benchmarks evaluate models primarily on English video-question pairs and generic content, without testing performance on non-English dialogues, region-specific contexts, or culturally diverse visual narratives. The absence of multilingual and multicultural evaluation for video LMMs forms a key motivation for ViMUL-Bench, which aims to fill that gap by assessing video understanding across diverse languages and cultural settings.
@@ 6 Conclusion
Acknowledgement: The computations were enabled through resources provided by NAISS at Alvis, partially funded by the Swedish Research Council under grant agreement no. 2022-06725, LUMI, hosted by CSC (Finland) and the LUMI consortium, and the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC. This work was partially supported by the Swedish Research Council (2022-04266), from KAW (DarkTree project; 2024.0076) and VR starting grant (2016-05543).
@@ Appendix I Performance based upon frames
      • Correct_Translated_MCQ_Choice_1, Correct_Translated_MCQ_Choice_2, Correct_Translated_MCQ_Choice_3, Correct_Translated_MCQ_Choice_4

### v3

@@ Abstract
Large multimodal models (LMMs) have recently gained attention due to their effectiveness to understand and generate descriptions of visual content. Most existing LMMs are in English language. While few recent works explore multilingual image LMMs, to the best of our knowledge, moving beyond the English language for cultural and linguistic inclusivity is yet to be investigated in the context of video LMMs. In pursuit of more inclusive video LMMs, we introduce a multilingual Video LMM benchmark, named ViMUL-Bench, to evaluate Video LMMs across 14 languages, including both low- and high-resource languages: Arabic, Bengali, Chinese, English, French, German, Hindi, Japanese, Russian, Sinhala, Spanish, Swedish, Tamil, and Urdu. Our ViMUL-Bench is designed to rigorously test video LMMs across 15 categories including eight culturally diverse categories, ranging from lifestyles and festivals to foods and rituals and from local landmarks to prominent cultural personalities. ViMUL-Bench comprises both open-ended (short and long-form) and multiple-choice questions spanning various video durations (short, medium, and long) with 8k samples that are manually verified by native language speakers. In addition, we also introduce a machine translated multilingual video training set comprising 1.2 million samples and develop a simple multilingual video LMM, named ViMUL, that is shown to provide a better tradeoff between high-and low-resource languages for video understanding. We hope our ViMUL-Bench and multilingual video LMM along with a large-scale multilingual video training set will help ease future research in developing cultural and linguistic inclusive multilingual video LMMs. Our proposed benchmark, video LMM codebase and training data are available at https://mbzuai-oryx.github.io/ViMUL/.
@@ A Culturally-diverse Multilingual Multimodal Video Benchmark & Model
Bhuiyan Sanjid Shafique1thanks: Equal contribution, Ashmal Vayani211footnotemark: 1, Muhammad Maaz1, Hanoona Abdul Rasheed1
Dinura Dissanayake1, Mohammed Irfan Kurpath1, Yahya Hmaiti2, Go Inoue1, Jean Lahoud1
Md. Safirur Rashid3, Shadid Intisar Quasem3, Maheen Fatima4, Franco Vidal2, Mykola Maslych2, Ketan Pravin More1, Sanoojan Baliah1, Hasindri Watawana1, Yuhao Li1, Fabian Farestam5, Leon Schaller6, Roman Tymtsiv7, Simon Weber6, Hisham Cholakkal1, Ivan Laptev1
Shin’ichi Satoh8, Michael Felsberg10 Mubarak Shah2, Salman Khan1,9, Fahad Shahbaz Khan1,10
1Mohamed bin Zayed University of Artificial Intelligence, 2University of Central Florida, 3Islamic University of Technology
4Air University, 5ETH Zurich, 6 Technische Universität München, 7Independent Researcher, 8National Institute of Informatics
9Australian National University, 10Linköping University
{bhuiyan.shafique, muhammad.maaz, hanoona.bangalath, dinura.dissanayake, mohammedirfan.k, go.inoue, jean.lahoud
ketan.more, sanoojan.baliah, hasindri.watawana, hisham.cholakkal, salman.khan, fahad.khan}@mbzuai.ac.ae
{ashmal.vayani, yohan.hmaiti, fr543419, mykola.maslych}@ucf.edu, {safirurrashid, shadidintisar}@iut-dhaka.edu
{leon.schaller, simon1.weber}@tum.de, {tymtsiv.roman, lyh88524}@gmail.com, maheen.fatima@students.au.edu.pk
ffarestam@student.ethz.ch, satoh@nii.ac.jp, michael.felsberg@liu.se, ivan.laptev@inria.fr, mubarak.shah@crcv.ucf.edu
@@ 1 Introduction
Existing efforts to explore linguistically and culturally diverse LMM benchmarks are limited to images Romero et al. (2024); Vayani et al. (2025). To the best of our knowledge, linguistic and cultural diversity are yet to be investigated for video LMMs. Video domain poses different challenges as it often depicts complex, culturally rich scenarios-local festivals, foods, rituals, or landmarks, that require understanding both the visual context and the language-specific narration or questions Swetha et al. (2025a, b); Kim et al. (2025); Sirnam et al. (2024). Fig. 1 illustrates an example where LMMs are asked to describe the taste of the Bengali dish Beef Tehari. The models fail to interpret the question in the local language, responding incorrectly and missing linguistic nuances. While short and long video understanding LMM benchmarks exist in literature, they are typically restricted to only English language. For instance, Video-MME Fu et al. (2024) focuses on diverse video analysis but in a single language, and MVBench Li et al. (2024b) emphasizes temporal reasoning (action sequences, motion) without multilingual considerations. Other recent efforts like ViLMA Kesen et al. (2024) and SEED-Bench Li et al. (2023b, a) probe video-language model’s abilities in zero-shot temporal grounding and procedural understanding, among other skills, yet none assess cross-lingual or cross-cultural comprehension. In short, there is no comprehensive benchmark to evaluate how well video LMMs perform across different languages and cultural contexts (see Tab. 1).
@@
  • We introduce ViMUL-Bench, a comprehensive benchmark for video LMMs covering 14 languages (including several under-represented ones) and 15 domains, including real-world cultural aspects. To our knowledge, this is the first effort to enable rigorous testing of video LMMs across a wide linguistic and cultural spectrum, emphasizing both cross-lingual and cultural comprehension (see Tab. 1).
@@ 2 Related Works
Table 1: Comparison of video LMM benchmarks emphasizing multilingual and cultural understanding. Domains represent the aspects covered by each dataset for different languages. Annotation Type is categorized as follows: Human - Questions were created in the local language. Human+Auto - Questions were generated or translated using GPT-4/Google API and later validated by human experts. Auto: Questions were generated or translated automatically without human validation. ‘-’ indicates that information is not available.
@@
Video LMM Benchmarks: Extending multimodal evaluation to video introduces additional challenges due to temporal dynamics Swetha et al. (2021, 2023). A number of video-language benchmarks have emerged to assess LMMs on video understanding, though they focus on general monolingual capabilities. Video-MME Fu et al. (2024) is introduced as the first comprehensive evaluation of LMMs on video analysis, covering a broad spectrum of video domains. It offers a full-spectrum evaluation, covering questions from perception to reasoning, and variable videos lengths ($\sim$10 secs to $\sim$1 hour). Another effort, MVBench Li et al. (2024b) concentrates on temporal reasoning skills and proposes an automatic pipeline to generate a large multiple-choice QA benchmark by leveraging existing video datasets and GPT-based annotators. Beyond general benchmarks, ViLMA Kesen et al. (2024) takes a fine-grained approach as it uses carefully designed counterfactual video pairs to probe a model’s temporal grounding and linguistic understanding in a zero-shot setting. SEED-Bench is another comprehensive multimodal benchmark that includes some video-based questions; however, those are largely confined to temporal or procedural understanding tasks Li et al. (2023b, a). These benchmarks have a significantly advanced evaluation for video-based multimodal reasoning and generation, however they lack multilingual and culturally aware components. Current video benchmarks evaluate models primarily on English video-question pairs and generic content, without testing performance on non-English dialogues, region-specific contexts, or culturally diverse visual narratives. The absence of multilingual and multicultural evaluation for video LMMs forms a key motivation for ViMUL-Bench, which aims to fill that gap by assessing video understanding across diverse languages and cultural settings.
@@ 6 Conclusion
Acknowledgement
The computations were enabled through resources provided by NAISS at Alvis, partially funded by the Swedish Research Council under grant agreement no. 2022-06725, LUMI, hosted by CSC (Finland) and the LUMI consortium, and the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC. This work was partially supported by the Swedish Research Council (2022-04266), from KAW (DarkTree project; 2024.0076) and VR starting grant (2016-05543).
@@ Appendix I Performance based upon frames
      • Correct_Translated_MCQ_Choice_1
      • Correct_Translated_MCQ_Choice_2
      • Correct_Translated_MCQ_Choice_3
      • Correct_Translated_MCQ_Choice_4